### v2

# Angular momentum of rotating fermionic superfluids by Sagnac phonon interferometry

Abstract
Fermionic many-body systems provide an unrivaled arena to investigate how interactions drive the emergence of collective quantum behavior, such as macroscopic coherence and superfluidity 1. Central to these phenomena is the formation of Cooper pairs, correlated states of two fermions that behave as composite bosons and condense below a critical temperature. However, unlike elementary bosons, these pairs retain their internal structure set by underlying fermionic correlations, essential for understanding superfluid properties throughout the so-called Bose–Einstein condensate (BEC) to Bardeen–Cooper–Schrieffer (BCS) crossover — a cornerstone of strongly correlated fermionic matter 2. Here, we harness a sonic analog of the optical Sagnac effect 3 to disclose the composite nature of fermionic condensates across the BEC–BCS crossover. We realize an in-situ loop interferometer by coherently exciting two counter-propagating long-wavelength phonons of an annular fermionic superfluid with tuneable interparticle interactions. The frequency degeneracy between clock- and anticlock-wise sound modes is lifted upon controllably injecting a quantized supercurrent in the superfluid ring 4, resulting in a measurable Doppler shift that enables us to probe the elementary quantum of circulation and the angular momentum carried by each particle in the fermionic fluid 5, 6. Our observations directly reveal that the superflow circulation is quantized in terms of $h/2m$, where $m$ is the mass of the constituents, in striking contrast to bosonic condensates where $h/m$ is the relevant circulation quantum. Further, by operating our interferometer at tunable temperature, we measure the thermal depletion of the superfluid in the unitary Fermi gas, demonstrating phonon interferometry as a powerful technique for probing fundamental properties of strongly-correlated quantum systems.
Rotating a quantum system is amongst the most effective strategies for both uncovering its macroscopic behavior and linking it to microscopic features. A hallmark of superfluidity, as anticipated by Onsager 7 and Feynman 8, is the quantization of circulation that emerges from the single-valuedness of the macroscopic wavefunction describing the entire quantum system (see Ref.[ 9] for a recent review). Superfluid flow is constrained by the discrete quantization of its circulation $\oint{\bf v}_{s}\cdot\mathrm{d}{\bf l}$ into units of $\kappa$, i.e. the quantum of circulation, which provides insight into the very mechanisms driving the emergence of superfluidity. Pioneering measurements of $\kappa$ in superfluid 3He-B 10, 11, 12 yielded $h/2m$ (where $h$ is Planck’s constant and $m$ mass of a single atom), contrasting with He-II result of $h/m$ 13, 14, 15 and thereby confirming that transport in 3He occurs through Cooper pairs of opposite spin rather than single particles 16. Similarly, for superconductors the single-valuedness of the electronic wavefunction enforces flux quantization, $\Phi_{0}=\oint{\bf A}\!\cdot\!\mathrm{d}{\bf l}$, with the vector potential $\bf A$} playing the role of the superfluid velocity in the case of charged particles. Early observations in superconducting rings confirmed charge-$2e$ Cooper pairs 17, 18 for standard superconductors, and recent studies suggest charge-$4e$ and $6e$ states 19 for more exotic ones. In ultracold atomic Fermi gases, the formation of Abrikosov vortex lattices in rotating traps offered the first striking evidence of superfluidity across the BEC–BCS crossover 20. However, only indirect estimates of $\kappa$ could be extracted there from the inter-vortex spacing at a given rotation velocity. Complementary insights came from angular momentum measurements in slowly rotating Fermi gases, which revealed deviations of the moment of inertia from the classical rigid-body value highlighting the role of temperature 21, 22.
In this work, we realize a Sagnac-like phonon interferometer to measure the quantum of circulation $\kappa$ in annular atomic Fermi superfluids across the BEC–BCS crossover. In analogy to optical Sagnac gyroscopes that detect rotation-induced phase shifts between counter-propagating light beams, we exploit a rotation-induced phononic phase shift to precisely quantify the angular momentum per particle. We find $\kappa$ to equal $h/2m$ within experimental uncertainties, both in the BEC regime of tightly bound dimers as well as in the strongly interacting unitary Fermi gas of extended pairs. We thus realize the first direct, in-situ rotation sensing protocol for fermionic quantum fluids, complementing previous measurements in rotating bosonic condensates 23, where $\kappa$ was found to equal $h/m$. Leveraging the periodic boundary conditions of an annular trapping geometry, we excite two long-lived counter-propagating phonon modes $(\omega_{\pm},\pm\textbf{k})$, whose interference creates a circular standing-wave density pattern with periodicity $2\pi/k$. Circulating currents alter the propagation velocity $c_{s}$ of these modes owing to a Doppler shift $\Delta_{c}$, effectively modifying the sonic dispersion relation as $\omega_{\pm}=c_{s}k\pm\Delta_{c}k$. This frequency splitting leads to a precession of the standing wave at angular frequency $\Omega=\Delta_{c}k$, such that the enclosed Sagnac phase appears as the angular shift $\Delta\theta=\Omega t$ (see Fig. 1a-b).
Figure 1: Sagnac-like matter-wave interferometer with annular Fermi superfluids. In analogy with optical Sagnac interferometers utilizing counter-propagating light waves in a loop, our interferometer is based on the excitation of two counter-propagating sound modes in an annular fermionic-pair condensate with tunable interactions. In the absence of rotation (a), both excitations follow the dispersion relation $\omega_{0}=c_{s}k$. In a rotating system (b), they experience a relative Doppler shift, yielding $\omega_{\pm}=c_{s}k\pm\Omega$, where the frequency splitting $\Omega$ is proportional to the angular momentum per particle. Such Doppler splitting manifests as a precession of the standing-wave atomic density pattern, accruing a phase shift $\Delta\theta=\Omega t$ over a time $t$. (c), In our phonon loop-interferometer, the longest-wavelength phonon modes with $\textbf{k}\bar{R}=\pm\hat{\theta}$ are simultaneously excited by a weak potential perturbation $V_{\text{pert}}=-V_{0}\cos{\theta}$ that is suddenly quenched off at $t=0$, defining the onset of the dynamics. (d), In-situ atomic density profiles recorded at different evolution times reveal the precession of the standing wave, with the angular shift $\Delta\theta=\Omega t$, from which we experimentally extract $\Omega$. Here, $\tau=2\pi/\omega_{0}$ denotes the phonon oscillation period.
At zero temperature, the Doppler shift $\Delta_{c}$ coincides with the background superfluid velocity $v_{s}$, as expected from a Galilean transformation that brings the system into the rest frame of the superfluid. The identity $\Delta_{c}=v_{s}$ has profound implications, as $v_{s}$ is quantized and proportional to $\kappa$, according to $v_{s}=\textit{w}\kappa/(2\pi\bar{R})$, where $\textit{w}=0,\pm 1,\pm 2,\dots$ denotes the integer winding number of the persistent current, and $\bar{R}$ the mean ring trap radius. Consequently, measuring $\Delta_{c}$ would allow to directly probe that: (i) the quantum of circulation for a Fermi superfluid is $\kappa=h/2m$ regardless of the interaction regime, and (ii) the angular momentum per particle $\ell_{z}=m\bar{R}v_{s}$ is quantized according to the law $\ell_{z}=m\textit{w}\kappa/2\pi$.
At finite temperature, Landau’s two-fluid hydrodynamic theory 24 predicts the propagation of two distinct sound modes, known as first and second sound. When the velocities of the normal and superfluid components are different, the Doppler shifts of the two sounds exhibit a more complex dependence 25, 26, 27, 28. As first pointed out by Khalatnikov 29, in a weakly compressible fluid such as liquid 4He, the Doppler shift of first sound—being a density wave—is fixed by the fluid flow according to the relation $\Delta_{c}=f_{s}v_{s}+f_{n}v_{n}$, where $v_{s,n}$ are the velocities of the superfluid and normal components, and $f_{s,n}=N_{s,n}/N$ are their respective fractions, with $f_{s}+f_{n}=1$. Note that only if $v_{n}=v_{s}$ the Doppler shift reduces to the velocity of the fluid. In the same limit of weakly compressible fluids, the Doppler effect of second sound exhibits a different anomalous behavior 29. The above relation for $\Delta_{c}$ is expected to hold also in superfluid Fermi gases near unitarity (see Methods). An interesting situation arises when the circulating current is carried solely by the superfluid component, i.e., when $v_{n}=0$. In this case, the measurement of the Doppler shift $\Delta_{c}$ provides access to $\ell_{z}$ according to the relation $\ell_{z}=m\bar{R}\Delta_{c}$, taking the form
$\ell_{z}=\frac{m\textit{w}}{2\pi}\,\frac{N_{s}}{N}\;\kappa,$ (1)
which reduces to the $T=0$ case for $N=N_{s}$. The result in Eq. (1) opens the interesting possibility of measuring the superfluid density at finite temperature, as we will discuss later.
Figure 2: Phonon interferometry across the BEC–BCS crossover. Temporal evolution of the azimuthal density modulation $\delta n(\theta,t)$ for a BEC at $1/k_{F}a=3.31(7)$ (a) and a unitary Fermi gas (b), both in the absence of circulation. Dashed lines in panels a(i) and b(i) are a guide to the eye of the nodal points of the standing wave. Panels a(ii) and b(ii) display the corresponding density-weighted angular moments $\langle\cos\theta\rangle$ (brown) and $\langle\sin\theta\rangle$ (black), as defined in the main text. For visual clarity, $\langle\cos\theta\rangle$ is vertically offset by +0.1+0.1 and $\langle\sin\theta\rangle$ is multiplied by 3. To extract the phonon frequency $\omega_{0}$, we fit the density modulations $\delta n(\theta,t)$ using Eq. (2); the resulting fits are shown as continuous lines on the moments traces. (c), The phonon velocity, expressed as $c_{s}/v_{F}=\omega_{0}\bar{R}/v_{F}$ with $v_{F}=\sqrt{2mE_{F}}$ the Fermi velocity. Insets show the corresponding power spectral densities of panels a(i) and b(i), confirming the monochromatic excitation with $\bar{R}k=\pm 1$, and the absence of second-sound coupling. (d), Quality factor $Q=\omega_{0}/\Gamma$ across the BEC–BCS crossover. (e), Damping rate $\Gamma/k^{2}$ expressed in units of $\hbar/m$. Error bars in panels c-e stem from the standard fitting error of $\delta n$.
[…]
The interferometric sequence begins by raising a weak perturbation potential $V_{\text{pert}}=-V_{0}\cos(\theta)$ (see Fig. 1c). The perturbation strength $V_{0}$ is kept small to ensure operation within the linear-response regime (see Methods). We allow the system to equilibrate for 50 ms, damping the residual motion of the normal component 21 along with other spurious excitations. The perturbation is then suddenly turned off, generating two counter-propagating monochromatic excitations with the lowest available wavenumber in the system, $|\mathbf{k}|=1/\bar{R}$ (see Fig. 1d). We note that similar phonon interferometric schemes have been explored in annular bosonic superfluids 5, 31, 6.
Figure 3: Phonon Doppler shift and angular momentum per particle across the BEC–BCS crossover. Temporal evolution of the azimuthal density modulation $\delta n(\theta,t)$ in a rotating superfluid with a single quantum of circulation ($\textit{w}=1$) for a BEC at $1/k_{F}a=3.31(7)$ (a) and a unitary Fermi gas (b). The color-scales are the same of Fig. 2a-b, respectively. Dashed lines in panels a(i) and b(i) are guides to the eye of the nodal points of the standing wave. Panels a(ii) and b(ii) display the corresponding density-weighted angular moments $\langle\cos\theta\rangle$ (brown) and $\langle\sin\theta\rangle$ (black); for clarity, $\langle\cos\theta\rangle$ is offset by +0.1+0.1 and $\langle\sin\theta\rangle$ is multiplied by 3. The oscillatory behavior of $\langle\sin\theta\rangle(t)$ associated to the precession of the density modulation provides an unambiguous signature of rotation. We extract the Doppler shift $\Delta_{c}$, fitting $\delta n(\theta,t)$ with Eq. (2), as a function of the total imprinted phase employed to generate the persistent current for the BEC (c), and UFG (d) regimes. In addition, we show the mean winding number $\langle\textit{w}\rangle$ measured over repeated acquisitions after the imprinting protocol. (e), Measured angular momentum per particle $\ell_{z}=m\bar{R}\Delta_{c}$ as a function of winding number w for a BEC at $1/k_{F}a=3.31(7)$ (green squares), a unitary Fermi gas (blue circles), and a BCS superfluid at $1/k_{F}a=-0.55(1)$ (orange triangles). (f), Angular momentum per particle across the BEC–BCS crossover. The black cross marks the critical interaction strength below which persistent currents are no longer observed. Colors of symbols identify measurements in panel e. The dashed line marks $\ell_{z}/\hbar\textit{w}=1/2$. Error bars in panels c-d stem from the standard fitting error of $\delta n$, whereas in panels e-f they represent the standard error of the mean over at least 3 repetitions.
We first test the interferometer response for non-rotating systems. For each interaction regime throughout the crossover, the time-evolution of the interference pattern is tracked by imaging the in-situ atomic density. From these measurements, we extract the normalized azimuthal density profile defined as $n_{\theta}(\theta)=\int n_{\text{3D}}(r,\theta,z)\,r\,dr\,dz/N$. Figure 2a-b shows the temporal evolution of the density modulation, $\delta n(\theta,t)=n_{\theta}(\theta,t)-n_{\theta}^{0}$, where the equilibrium profile $n_{\theta}^{0}$ is determined from measurements of the unperturbed cloud. At finite temperature, density excitations generally couple to both first and second sound modes, resulting in two characteristic modulation frequencies for each wavenumber $k$. The relative contribution of each mode to the density is governed by the density response function $\chi(k,\omega)$. Recent characterizations of $\chi(k,\omega)$ at unitarity indicate that for temperatures below $0.7\,T_{c}$, second sound couples only weakly to density 32, manifesting mostly as a entropy wave 33. In the temperature range explored here, our excitation protocol selectively excites only two counter-propagating first sound modes. This is confirmed by the power spectral density of the measured signal, which displays the peaks of the clockwise $(\pm\omega_{0}=\pm c_{s}\bar{R}^{-1})$ and anti-clockwise $(\pm\omega_{0}=\mp c_{s}\bar{R}^{-1})$ phonon modes (see insets of Fig. 2c), without populating additional peaks – e.g. at frequencies $|\omega|\approx 0.26\,\omega_{0}$, expected for second sound excitations 33, 32. In the absence of a circulating current, the density modulations resulting from quenching off the static perturbation $V_{\text{pert}}$ takes the form $\delta n(\theta,t)=e^{-\Gamma t/2}\,V_{0}(\chi/2)\cos(\theta)\cos(\omega_{0}t)$, with a time dependence fixed by the first sound frequency $\omega_{0}$ and the damping coefficient $\Gamma$, which sets the decay of the collective oscillations. Here, $\chi$ is the static response that coincides with the isothermal compressibility $(\partial\rho/\partial p)_{T}$ according to the compressibility sum rule 34. Furthermore, we analyze the density-weighted angular moments $\langle\cos{\theta}\rangle(t)=\int_{0}^{2\pi}n_{\theta}(\theta,t)\cos\theta\,d\theta$ and $\langle\sin{\theta}\rangle(t)=\int_{0}^{2\pi}n_{\theta}(\theta,t)\sin\theta\,d\theta$. In these observables, the presence of circulating currents is directly revealed by $\langle\sin{\theta}\rangle(t)\neq 0$. As shown in Fig. 2a-b, this is consistently vanishingly small in the absence of superflow around the ring. By using the above relations, we fit the evolution of $n_{\theta}(\theta,t)$ to extract the phonon propagation speed $c_{s}$ and the corresponding $\Gamma$. The extracted values of $c_{s}$, shown in Fig. 2c, increase monotonically from the BEC to the BCS regime, reflecting the rise in chemical potential through the crossover. Nevertheless, they are consistently lower than those found in homogeneous systems 35, primarily due to the vertical density inhomogeneity of our trapped gases 36. At unitarity, the sound velocity is directly connected to the Bertsch parameter $\xi_{B}$, through its link to the thermodynamic compressibility 34. At the lowest temperature explored in this work, ($T/T_{c}\sim 0.4$), we measure $\xi_{B}=$0.396(29)\text{\,}$\$ (see Methods). This value is consistent with previous determinations from thermodynamic measurements 37, 38, sound-excitation experiments 32, 39, and quantum Monte Carlo calculations 40, 41.
The interferograms display high contrast over several phonon rotation periods, reflecting the long-lived character of long-wavelength sound modes in the ring, especially in the strongly interacting regime. To characterize this behavior more quantitatively, we extract the interferometer quality factor $Q=\omega_{0}/\Gamma$, from the phonon damping rate. The results, shown in Fig. 2d for each interaction strength, reveal an enhancement of $Q\approx 30$ for fermionic superfluids near unitarity. We can link this behavior to sound transport properties in strongly interacting systems. In the hydrodynamic collisional regime, the sound diffusivity $D=\Gamma/k^{2}$, approaches the fundamental limit fixed by the ratio $\hbar/m$, signaling the onset of quantum-limited transport 42, 43. Rescaling $\Gamma$ by $k^{2}$ yields the behavior shown in Fig. 2(e), which is consistent with previous results 33, 44, 45, 46, 47. As interactions are tuned away from unitarity towards the BEC side, $\Gamma/k^{2}$ first increases with the mean free path up to $1/k_{F}a\sim 2$, then decreases again, likely reflecting the onset of a dimensional crossover (see Methods). While the microscopic mechanisms governing phonon damping across the crossover and their sensitivity to the effective dimensionality remain open challenges for future studies, our results establish strongly interacting Fermi superfluids as a promising platform for atom interferometry and sensing 6.
To access the angular momentum per particle $\ell_{z}$, we set the superfluid gas into rotation by using a phase-imprinting technique 4, known to provide enhanced stability and control compared with conventional stirring methods 48, 23, 20. We inject quantized circulation of winding number w by optically imprinting a phase $\Delta\phi_{I}$ on the ring. In the presence of the persistent current, we excite the two phonon branches and study the Doppler effect on this rotating system. Unlike the non-rotating case, the density modulation $\delta n(\theta,t)$ exhibits a clear spatio-temporal drift, signaling a precession of the interference pattern (see Fig. 3a-b for $\textit{w}=1$). Moreover, $\langle\sin\theta(t)\rangle$ shows clear oscillations with non-zero amplitude, revealing the presence of the background circulating flow. To quantify the precession rate, we fit the observed dynamics with the theoretical prediction of linear response theory:
$\delta n(\theta,t)=\frac{V_{0}}{2}e^{-\frac{\Gamma t}{2}}\left[\chi_{+}\cos(\theta-\omega_{+}t)+\chi_{-}\cos(\theta+\omega_{-}t)\right].$ (2)
This expression explicitly accounts for the different time-dependent evolution of the two Doppler shifted components of the first sound mode, resulting in the precession $d\theta/dt=\Omega=\Delta_{c}/\bar{R}$ visible as a shift of the local zeros of $\delta n(\theta,t)$ (see guidelines in Fig. 3a-b). In the above equation, $\chi_{\pm}=(\chi/2)(1\mp\Delta_{c}/c_{s})$ are the corresponding contributions to the static response $\chi_{+}+\chi_{-}=\chi_{1}\simeq\chi$, neglecting the contribution of second sound to the response.
In our system, the background rotational flow is expected to be sustained solely by the superfluid, as any residual incoherent motion of the normal component is damped during the equilibration period preceding the quench of $V_{\text{pert}}$, effectively yielding $f_{n}v_{n}\simeq 0$ when the potential is quenched. To corroborate that the contribution of normal fluid motion is negligible, we measured the Doppler shift $\Delta_{c}$ as a function of the imprinted phase 4 $\Delta\phi_{I}$ (see Fig. 3c-d) and found that, mirroring the measured mean winding number $\langle\textit{w}\rangle$, it exhibits a step-like behavior rather than the linear monotonous increase expected for a normal moment of inertia 22, 21. The visible deviation between the phase imprinted to obtain $\langle\textit{w}\rangle=1$ in a BEC or in a UFG is inherent to the optical phase-imprinting technique in fermionic gases 4, 49. For the following discussions, interferometric measurements were performed using the central value of each $\Delta\phi_{I}$ step to establish a well-defined circulation state, w. Fig. 3e shows the angular momentum per particle as a function of the winding number w, for different interaction regimes. As expected, $\ell_{z}$ increases linearly with w, or equivalently with $v_{s}$, with a slope close to the zero-temperature value $\hbar/2$. A deviation is observed in the BCS regime, as discussed next. In Fig. 3f, we present the slope $\ell_{z}/\textit{w}$ as a function of $1/k_{F}a$. In the BEC regime, $\ell_{z}/\textit{w}\approx\hbar/2$, consistent with superfluidity arising from tightly bound molecules. A comparable value is found at unitarity, indicating that the superflow is carried by fermion pairs rather than individual particles. In the BCS regime, however, $\ell_{z}/\textit{w}$ is strongly suppressed, which we attribute to the marked reduction of the superfluid fraction as $T_{c}/T_{F}$ rapidly decreases in this limit 50. These results are consistent with the critical value of $1/k_{F}a$ beyond which the phase-imprinting protocol fails to generate persistent currents, $1/k_{F}a\lesssim$-0.7(1)\text{\,}$\$, indicated by a cross in Fig. 3f, in agreement with the observed trend of $\ell_{z}$.
Figure 4: Superfluid fraction at unitarity. We estimate the superfluid fraction extracted from the angular momentum per particle, obtained from the relation $N_{s}/N=2m\bar{R}\Delta_{c}/\hbar$, as a function of the reduced temperature $T/T_{c}$, where $T_{c}\simeq 0.18\,T_{F}$ in our trap 30. Here, we used the fact that the quantum of circulation at unitarity is $\kappa=h/2m$. Color solid lines represent previous experimental extractions of the superfluid fraction in homogeneous unitary Fermi gases, adjusted to account for harmonic confinement along the $z$-axis: Ref.[ 32] (green), Ref.[ 33] (brown), and Ref.[ 51] (red). Similarly, we compare our results with predictions 52 from mLPDA mean-field description (black dashed line) and extended GMB approach (black continuous line). Error bars represent the standard fitting error of $\delta n$, and $T/T_{c}$.
Owing to the near-vanishing of $v_{n}\simeq 0$, as anticipated in Eq. (1), the gas superfluid fraction at finite temperature can be quantitatively inferred from the temperature dependence of $\ell_{z}$. Motivated by this, we performed measurements of $\ell_{z}$ at unitarity as a function of temperature in the range $0.4-0.6~T_{c}$. The results, shown in Fig. 4, reveal a decrease in the superfluid fraction as the temperature increases. However, similarly to the measurements of $c_{s}/v_{F}$, the obtained values are lower than those reported for homogeneous systems 51, 33, 32, likely due to the vertical density inhomogeneity. For a meaningful comparison, we extrapolate the superfluid fraction of Refs.[ 51, 33, 32] accounting for the effect of harmonic confinement along the $z$-axis through the equation of state at unitarity 37, and the local density approximation (see Methods). Applying the same considerations, in Fig. 4 we compare also with the theoretical predictions 52 based on the extended Gorkov–Melik–Barkhudarov (GBM) approach to the Fermi superfluid phase, and the modified Local Phase Density Approximation (mLPDA) mean-field approach. Taking the vertical harmonic confinement into account brings our results into good agreement with both experiment and theory, confirming that the reduction of $\ell_{z}$ at finite temperature at unitarity and in the BCS regime originates from finite temperature effects which diminish the superfluid fraction. Our findings provide a novel benchmark for the low-temperature properties of paired superfluids and establish the factor-of-two reduction in the quantum of circulation, $\kappa=h/2m$, as a general hallmark of Fermi superfluidity. We stress that our interferometric technique measures the angular momentum per particle without relying on the precise knowledge of the system’s equation of state; it could thus be readily extended to diverse systems, ranging from two-dimensional 53, disordered 54 or periodically modulated 55, 56 superfluids to dipolar supersolids 57, 28, 58.
A further direction opened by our work is the investigation of the anomalous Doppler effect of second sound 29, 26, 27, 28 via thermographic techniques 33, thus allowing the study of heat and entropy transport in moving superfluids. To enable more direct comparisons with theoretical predictions and other experiments, it would be beneficial to employ 3D box-like trapping potentials that yield fully homogeneous atomic densities 59. The ability to generate sound excitations with distinct wavevectors using tailored optical perturbation potentials will allow also for the systematic study of sound damping mechanisms, such as phonon–phonon and phonon–single-particle interactions, across the BEC–BCS crossover. This approach offers a direct and effective route to investigate how the curvature of the phonon dispersion affects energy dissipation and its connection to the onset of wave turbulence in strongly interacting Fermi superfluids 35, 60, 61, 62.
Acknowledgements
We thank the Quantum Gases group at LENS for stimulating discussions. We also thank S. Giorgini, F. Dalfovo, A. Recati, M. Šindik and T. Zawiślak and G. Strinati for valuable comments about this work. We thank the European Center for Theoretical Studies in Nuclear Physics and Related Areas (ECT*) in Trento for support at the Workshop "Nonequilibrium phenomena in superfluid systems: atomic nuclei, liquid helium, ultracold gases, and neutron stars", favoring fruitful discussions with M. Zwierlein. G.R. and G.D.P. acknowledge financial support from the PNRR MUR project PE0000023-NQSTI. G.R. acknowledges funding from the Italian Ministry of University and Research under the PRIN2017 project CEnTraL and project CNR-FOE-LENS-2024. S.S. acknowledges support from the Provincia Autonoma di Trento. F.S. acknowledges support from the EU under the Horizon 2020 research and innovation program (project OrbiDynaMIQs, GA No. 949438). The authors acknowledge support from the European Union - NextGenerationEU within the “Integrated Infrastructure Initiative in Photonics and Quantum Sciences" (I-PHOQS). The authors acknowledge funding from INFN through the RELAQS project. This publication has received funding under the Horizon Europe programme HORIZON-CL4-2022-QUANTUM-02-SGA (project PASQuanS2.1, GA no. 101113690) and Horizon 2020 research and innovation programme (GA no. 871124).
Data availability
The data supporting the figures within this paper will be available in a Zenodo repository.
Author contributions
All authors contributed to the interpretation of the results and to the writing of the manuscript.
Competing interests
The authors declare no competing interests.